OpenStudy (anonymous):

Which of the following best describes how to use of the addition property of equality to isolate the variable x below?

OpenStudy (anonymous):

All that's going on is 5 is equal to x-2. Since they're both the same you can do the same thing to both sides and it won't matter! I'll show you what I mean. 3=3 right? So if you add 4 to both sides like this: 3+4=3+4 You can see that they just bring us to 7=7 Which makes perfect sense! Don't let the x throw you off! think of 5=x-2 like 3=3 in my example above. 5=x-2 So you want to find out what x is right? Well, you can just add 2 to both sides! 5+2=x-2+2 with me so far? Now you can see that -2 and +2 just makes 0 so 7=x See how we just used addition and kept both sides equal? That's all they mean by "addition property of equality!"

OpenStudy (anonymous):

thank u @Bri_Durbin

OpenStudy (anonymous):

no prob.